Now consider the same economy, and the central bank sells $8 billion worth of government
bonds to local banks. State the likely effects on the money supply under reserve requirements
of:
a) 10%.
b) 15%.
c) 20%.
d) 25%.
1
Expert's answer
2017-11-08T12:58:07-0500
a) 10%: the money supply will decrease by 8/0.1 = $80 billion.
b) 15%: the money supply will decrease by 8/0.15 = $53.33 billion.
c) 20%: the money supply will decrease by 8/0.2 = $40 billion.
d) 25%: the money supply will decrease by 8/0.25 = $32 billion.